# Bayesian optimisation of poloidal field coil positions in tokamaks

### v1

@@ 2.2 The Gaussian process surrogate
In BO, the most common type of surrogate used is a Gaussian process (GP), which is a probabilistic machine learning model for performing inference on the value of a function $f\colon\mathcal{U}\to\mathbb{R}$ given some training data—see Garnett (2023, Chp. 2). It is characterised by a mean function $m\colon\mathcal{U}\to\mathbb{R}$ and a positive semi-definite covariance function $k\colon\mathcal{U}\times\mathcal{U}\to\mathbb{R}$ (with $k(\cdot,\cdot)\geqslant 0$) such that the prior can be defined as
$f\sim\mathcal{GP}\left(m,k\right).$ (2.3)
@@
The distribution of this prior is the joint distribution of (infinitely) many Gaussian random variables and can be thought of as a distribution over functions. Therefore, at a finite set of evaluation points $\bm{X}=\{\bm{x}_{1},\bm{x}_{2},\ldots\}\subset\mathcal{U}$ we have that
$f(\bm{X})\sim\mathcal{N}\left(\bm{\mu},\bm{\Sigma}\right),$ (2.4)
@@
Training a GP requires conditioning the prior (2.3) on the dataset [^2] In practice, the data in $\mathcal{D}$ is transformed such that the outputs ($f(\bm{x})$) are standardised to have a mean of 00 and a variance of $1$. of known function evaluations $\mathcal{D}$ such that we obtain the following posterior distribution
$f(\bm{X})\mid\mathcal{D}\sim\mathcal{N}(\bm{\hat{\mu}},\bm{\hat{\Sigma}}).$ (2.5)
@@
The quality of this posterior distribution (in terms of the mean accuracy and variance calibration), however, is highly dependent on the choices made for the functions $m$ and $k$. A typical choice for the mean function is $m\equiv 0$, which assumes no prior knowledge of the function being modelled and ensures model predictions from (2.5) are heavily influenced by the training data. The choice of covariance kernel is formed via our prior belief in the expected behaviour of the true function being modelled (e.g. non-periodicity and smoothness). The covariance function used here is the Matérn-$(1/2)$ (or exponential) kernel
@@
where $\|\cdot\|_{2}$ denotes the Euclidean distance (Pandit and Infield, 2019). In addition, the parameters $\ell$ and $\sigma$ define the input length scale (smaller values produce more ‘wiggly’ functions) and the function noise (smaller values lead to lower predictive uncertainty in the function). The covariance kernel encodes the relationship between input points and the resulting covariance matrix quantifies how a change in one point influences changes in another across the domain. The hyperparameters $\ell$ and $\sigma$ are tuned (for example, using traditional non-Bayesian optimisation algorithms) to produce the best fit to the training data such that the marginal log-likelihood of the posterior (2.5) is maximised.
@@
2.3 The classifier
The aim of classification is to label each data point in the input space as either a failure (00) or non-failure ($1$). We model the probability of failure by ‘wrapping’ the GP surrogate model in a logistic function (Williams and Barber, 1998):
$p\left(\mathrm{non{\text{-}}failure}\ |\ \bm{x}\right)=\frac{1}{1+e^{-f(\bm{x})}}\quad\bm{x}\in\mathcal{U}.$
This transforms the GP from a regression model into a classifier model by squishing the GPs prediction into a probability. This probability can be converted into a binary prediction by wrapping it in a shifted Heaviside function (e.g returning $1$ if the probability is $\geqslant 0.5$, else 00).
The GP model with classification now has a Bernoulli likelihood $p(\mathcal{D}|f)$, making the calculation of the posterior distribution $p(f|\mathcal{D})$ analytically intractable—unlike in (2.5) where the prior, likelihood, and, therefore, posterior were all Gaussian (see Williams and Rasmussen (2006, Chp. 3.4)). To address this, we approximate the posterior using a variational distribution $q(f|\mathcal{D};\bm{\lambda})$, chosen such that its likelihood $q(D|f;\mathbf{\bm{\lambda}})$ is Gaussian—with $\bm{\lambda}$ parameterising the new distribution (Tran et al., 2015). The parameters $\bm{\lambda}$ are found by maximising the evidence lower bound
$\mathcal{L}(\bm{\lambda})\coloneqq\mathbb{E}_{q(D|f;\bm{\lambda})}\left[\log p(D|f)\right]-\mathrm{KL}\left[q(f;\bm{\lambda})\ \|\ p(f)\right].$
The first term represents the expected log likelihood (observing the training data given the probability distribution over functions), while the second term denotes the (non-negative) Kullback–Leibler divergence between the two distributions. Clearly if the KL divergence was zero (the distributions were identical), we would be maximising over the original (log) Bernoulli likelihood. Once the $\bm{\lambda}$ are found, the GP can be conditioned on the data (as was shown in Section 2.2) using the new Gaussian likelihood $q(D|f;\mathbf{\lambda})$.
@@ 2.4 The acquisition function
Based on knowledge from the trained GP and classifier, the acquisition function provides us with a way to estimate the informativeness of evaluating $\bm{f}$ at previously unseen values of $\bm{x}\in\mathcal{U}$. Depending on the task at hand, there are many possible choices of acquisition functions, each tailored to specific objectives. As mentioned before, the key factor in selecting an appropriate one is that it should be computationally cheap (compared to $\bm{f}$) to evaluate given the surrogate model.
@@
Here, we use the expected hypervolume improvement (EHVI) function, which seeks to quantify the expected improvement in the hypervolume of $\mathcal{P}_{\bm{f}}$ when adding a new point to the dataset $\mathcal{D}$ (Daulton et al., 2020). EHVI is particularly suited to multi-objective optimization, as it effectively balances exploration and exploitation by focusing on regions of the search space that are both uncertain and potentially optimal. The EHVI function $\alpha_{\text{EHVI}}\colon\mathcal{U}\to\mathbb{R}$ is given by
$\alpha_{\text{EHVI}}(\bm{x})=\mathbb{E}_{\bm{f}}\left[\text{HV}(\mathcal{P}_{\bm{f}}\cup\{\bm{f}(\bm{x})\})-\text{HV}(\mathcal{P}_{\bm{f}})\right],$
where $\text{HV}(\mathcal{P}_{\bm{f}})$ is the hypervolume of the Pareto front (relative to some reference point [^3] The reference point is a fixed location in the objective space that does not belong to $\mathcal{P}_{\bm{f}}$ and ensures the hypervolume calculation does not diverge to infinity (Yang et al., 2019).) and $\mathbb{E}_{\bm{f}}$ is the expectation operator of (2.5) (with respect to the $l$ objectives, not the constraints).
The goal is to use constrained single-objective (non-Bayesian) gradient-based optimisation [^4] For example, a number of different BO packages use the Broyden–Fletcher–Goldfarb–Shanno optimisation algorithm (Byrd et al., 1995; Picheny et al., 2023). to find
$\bm{x}^{*}=\underset{\bm{x}\in\mathcal{U}}{\arg\max}\ \alpha_{\text{EHVI}}(\bm{x}).$
Once found, this point can be evaluated using $\bm{f}$ and added to the dataset $\mathcal{D}$.
@@
As mentioned before, we have both constraints on the function $\bm{f}$ and failure regions in the input space (that we wish to avoid). To do this, we can extend EHVI by multiplying it by a feasibility indicator $\Delta(\bm{x})\in\{0,1\}$, such that we obtain the Expected Constrained Hypervolume Improvement (ECHVI) function
@@
$\alpha_{\text{ECHVI}}(\bm{x})=\Delta(\bm{x})\alpha_{\text{EHVI}}(\bm{x}).$
The feasibility indicator $\Delta(\bm{x})$ returns $1$ if the probability of feasibility is above some threshold in $[0,1]$, otherwise it returns 00 (Gardner et al., 2014). The probability of feasibility is low if a point $\bm{x}$ is likely to violate any of the $m$ constraints or reside in a failure region. This means that a point that is Pareto optimal but likely infeasible would yield an ECHVI value of zero and would therefore not maximise the acquisition function.
3 The poloidal field coil design problem
@@
3.1 The STEP baseline design
We will be working with the initial PF coil setup and limiter geometry from the baseline STEP design presented by Hudoba et al. (2023a). The design is shown in Figure 3.1 and the information available to us from the baseline dataset are the:
@@ 3.1 The STEP baseline design
    permissible zones for each PF coil (i.e. the region of the $RZ$-plane in which each coil can be placed without intersecting the TF coils, diagnostic systems, or other parts of the tokamak).
@@
    limiter contour that confines the plasma equilibrium [^5] This was constructed using the strike plate locations and made to match the geometry illustrated in Tholerus et al. (2024)..
@@
    strike plate locations (i.e. segments of the limiter in the inner and outer divertor where the legs of the plasma separatrix will strike).
@@
Figure 3.1: The baseline setup from Hudoba et al. (2023a) in the $RZ$-plane (due to vertical symmetry about $Z=0$ only the top half of the tokamak is shown). The separatrix (red) outlines the shape of the plasma core and the divertor legs which hit the inner (green) and outer (blue) strike plates. The initial location of the seven PF coils (grey rectangles) are displayed within their respective permissible zones (red rectangles). Note the absence of the central solenoid, which is not used in the flat-top phase of operation shown here. The isoflux constraints (black crosses) define locations at which the separatrix should pass through.
3.2 Optimisation problem
@@
3.2.1 Input space
As can be seen in Figure 3.1, there are seven up-down symmetric [^6] Note that these coils are up-down symmetric around $Z=0$ and so when the upper coil is moved to some location $(R^{c},Z^{c})$ for $Z>0$ (i.e. in Figure 3.1), its corresponding lower coil is moved to the vertically opposite location $(R^{c},-Z^{c})$. PF coil circuits each with their own $(R^{c},Z^{c})$ centroid coordinate that is allowed to move freely such that no part of the coil leaves the permissible zone. The exceptions are the two QS coils which are only able to move vertically. This results in a twelve dimensional input space for the optimisation problem: five pairs of $(R^{c},Z^{c})$ coordinates for the PF coils and one $Z^{c}$ coordinate for each of the two QS coils. We normalise each of the coordinates with respect to their own permissible zones so that we can work with the unit hypercube $[0,1]^{12}$ as our input space. A more detailed explanation of the normalisation process can be found in Appendix A.
3.2.2 Objectives
@@
Figure 3.2: The baseline setup from Figure 3.1, this time marked with objective and constraint quantities. Shown are the separatrix (light pink), the upper X-point (red cross), the O-point/magnetic axis (green circle), and the inner (green star) and outer (red star) strikepoints. Also shown are the flux surfaces traced out to calculate the ICL (blue plus red lines) and OCL (yellow plus green lines).
@@
The second objective is to maximise the average of the inner connection length (ICL) and the outer connection length (OCL). The connection length is the distance traced out by a helical (i.e. moving both poloidally and toroidally) magnetic field line that starts at the inner (outer) edge of the last closed flux surface (LCFS) at the midplane and ends at the inner (outer) strikepoint (Doyle et al., 2021; Albanese et al., 2023). We refer to these inner and outer midplane points as the IMP and OMP, respectively. Larger connection lengths ensure that hot plasma leaving the core edge region will travel a further distance and therefore cool to more acceptable temperatures before hitting the strike plate. This is an important aspect of managing heat loads in the divertor region.
@@
The ICL and OCL are calculated using a 3D field line tracing algorithm, i.e. by carrying out an integral over the length of the magnetic field line of interest (Kos et al., 2019). By parametrising the position vector $\bm{r}(\ell)=\left(R(\ell),\phi(\ell),Z(\ell)\right)$ in terms of a length $\ell$, we can find the connection length $L$ by first solving the ordinary differential equations
$\frac{\mathrm{d}\bm{r}(\ell)}{\mathrm{d}\ell}=\bm{B}\left(\bm{r}(\ell)\right)=\begin{cases}-\frac{1}{R}\frac{\partial\psi(R,Z)}{\partial Z},\\[6.0pt] \frac{F(\psi(R,Z))}{R},\\[6.0pt] \frac{1}{R}\frac{\partial\psi(R,Z)}{\partial R},\end{cases}\quad\ell\in[0,L],$ (3.2)
@@
with an initial condition $\bm{r}(0)=\left(R(0),\phi(0),Z(0)\right)$. Here, $\bm{B}$ denotes the magnetic field, $\psi(R,Z)$ the scalar poloidal magnetic flux, and $F$ the toroidal magnetic field profile. To do this, we use a fourth-order Runge-Kutta method (with appropriately chosen step size $\Delta\ell$) and integrate until some terminating condition is met (see below), recording the value of $L$ obtained (i.e. the total number of steps $\Delta\ell$).
@@
In practice, however, the calculation is split into two stages. The first stage involves integrating from an initial location $3$mm radially outside the IMP/OMP and ending at some very small distance away from the upper X-point. The second stage integrates between the inner/outer strikepoint and the point close to the upper X-point. The length of these individual sections is then combined to return the final connection length. The two flux surfaces traced out when calculating the ICL and OCL are visualised in Figure 3.2. These techniques reduce the likelihood of the integrator getting stuck at the exact X-point, travelling around the LCFS (instead of going up into a divertor), and from travelling into the wrong divertor.
3.2.3 Constraints

### v2

@@ II.2 The Gaussian process surrogate
In BO, the most common type of surrogate used is a Gaussian process (GP), which is a probabilistic machine learning model for performing inference on the value of a function $f\colon\mathscr{U}\to\mathbb{R}$ given some training data—see Garnett (2023, Chp. 2). It is characterised by a mean function $m\colon\mathscr{U}\to\mathbb{R}$ and a positive semi-definite covariance function $k\colon\mathscr{U}\times\mathscr{U}\to\mathbb{R}$ (with $k(\cdot,\cdot)\geq 0$) such that the prior can be defined as
$f\sim\mathscr{GP}\left(m,k\right).$ (3)
@@
The distribution of this prior is the joint distribution of (infinitely) many Gaussian random variables and can be thought of as a distribution over functions. Therefore, at a finite set of evaluation points $\bm{X}=\{\bm{x}_{1},\bm{x}_{2},\ldots\}\subset\mathscr{U}$ we have that
$f(\bm{X})\sim\mathscr{N}\left(\bm{\mu},\bm{\Sigma}\right),$ (4)
@@
Training a GP requires conditioning the prior (3) on the dataset of known function evaluations $\mathscr{D}$ (with outputs standardised to mean 00, standard deviation $1$ and the inputs transformed to the unit hypercube) such that we obtain the following posterior distribution
$f(\bm{X})\mid\mathscr{D}\sim\mathscr{N}(\bm{\hat{\mu}},\bm{\hat{\Sigma}}).$ (5)
@@
The quality of this posterior distribution (in terms of the mean accuracy and variance calibration), however, is highly dependent on the choices made for the functions $m$ and $k$. A typical choice for the mean function is $m\equiv 0$, which assumes no prior knowledge of the function being modelled and ensures model predictions from (5) are heavily influenced by the training data. The choice of covariance kernel is formed via our prior belief in the expected behaviour of the true function being modelled (e.g. non-periodicity and smoothness). The covariance function used here is the Matérn-$(1/2)$ (or exponential) kernel
@@
where $\|\cdot\|_{2}$ denotes the Euclidean distance (Pandit and Infield, 2019). In addition, the parameters $\ell$ and $\sigma$ define the input length scale (smaller values produce more ‘wiggly’ functions) and the function noise (smaller values lead to lower predictive uncertainty in the function). The covariance kernel encodes the relationship between input points and the resulting covariance matrix quantifies how a change in one point influences changes in another across the domain. The hyperparameters $\ell$ and $\sigma$ are tuned (for example, using traditional non-Bayesian optimisation algorithms) to produce the best fit to the training data such that the marginal log-likelihood of the posterior (5) is maximised.
@@
II.3 The classifier
The aim of binary classification is to label each data point in the input space as either a failure (00) or a non-failure ($1$). The probability that a point $\bm{x}\in\mathscr{U}$ is a non-failure is modelled using a GP over a latent function, which is then transformed via the logistic function(Williams and Barber, 1998):
$p\left(\mathrm{non{\text{-}}failure}\ |\ \bm{x}\right)=\frac{1}{1+e^{-f(\bm{x})}}\quad\bm{x}\in\mathscr{U}.$ (6)
Using the logistic function transforms the GP from a regression model (as it is used in Section II.2) to a classifier by mapping its prediction into a probability that a point belongs to the non-failure class.
The GP model with classification now has a Bernoulli likelihood $p(\mathscr{D}|f)$, making the calculation of the posterior distribution $p(f|\mathscr{D})$ analytically intractable—unlike in (5) where the prior, likelihood, and, therefore, posterior were all Gaussian (see Williams and Rasmussen (2006, Chp. 3.4)). To address this, we approximate the posterior using a variational distribution $q(f|\mathscr{D};\bm{\lambda})$, chosen such that its likelihood $q(\mathscr{D}|f;\mathbf{\bm{\lambda}})$ is Gaussian—with $\bm{\lambda}$ parameterising the new distribution (Tran, Ranganath, and Blei, 2015). The parameters $\bm{\lambda}$ are found by maximising the evidence lower bound
$\mathscr{L}(\bm{\lambda})\coloneqq\mathbb{E}_{q(\mathscr{D}|f;\bm{\lambda})}\left[\log p(\mathscr{D}|f)\right]-\mathrm{KL}\left[q(f;\bm{\lambda})\ \|\ p(f)\right].$
The first term represents the expected log likelihood (observing the training data given the probability distribution over functions), while the second term denotes the (non-negative) Kullback–Leibler divergence between the two distributions. Clearly if the KL divergence was zero (the distributions were identical), we would be maximising over the original (log) Bernoulli likelihood. Once the $\bm{\lambda}$ are found, the GP can be conditioned on the data (as was shown in Section II.2) using the new Gaussian likelihood $q(\mathscr{D}|f;\mathbf{\lambda})$.
@@ II.4 The acquisition function
Based on knowledge from the trained GP and classifier, the acquisition function provides us with a way to estimate how informative evaluating $\bm{f}$ at a previously unseen point $\bm{x}\in\mathscr{U}$ will be. Depending on the task at hand, there are many possible choices of acquisition functions, each tailored to specific objectives. As mentioned before, the key factor in selecting an appropriate one is that it should be computationally cheap (compared to $\bm{f}$) to evaluate given the surrogate model.
Here, we use the expected hypervolume improvement (EHVI) function, which seeks to quantify the expected increase in the hypervolume of $\mathscr{P}_{\bm{f}}$ when adding a new point to the dataset $\mathscr{D}$ (Daulton, Balandat, and Bakshy, 2020). The hypervolume HV of $\mathscr{P}_{\bm{f}}$ is defined as the $l$-dimensional integral of the subspace
$\{\bm{y}\in\mathbb{R}^{l}\mid\exists\bm{p}\in\mathscr{P}_{\bm{f}}\ \text{s.t.}\ \bm{p}\prec\bm{y}\},$
dominated by $\mathscr{P}_{\bm{f}}$ (Yang et al., 2019).
@@
EHVI is particularly suited to multi-objective optimisation, as it effectively balances exploration and exploitation by focusing on regions of the search space that are both uncertain and potentially optimal. The EHVI function $\alpha_{\text{EHVI}}\colon\mathscr{U}\to\mathbb{R}$ is given by
$\alpha_{\text{EHVI}}(\bm{x})=\mathbb{E}_{\bm{f}}\left[\text{HV}(\mathscr{P}_{\bm{f}}\cup\{\bm{f}(\bm{x})\})-\text{HV}(\mathscr{P}_{\bm{f}})\right],$
where $\mathbb{E}_{\bm{f}}$ is the expectation operator of (5) (with respect to the $l$ objectives, not the constraints). Recalling that $\bm{f}(\bm{x})$ is a random variable, this function describes how much additional volume in objective space we expect to gain by sampling at a new point $\bm{x}$, relative to the current Pareto front. Please refer to Yang et al. (2019) for a more rigorous treatment of this material.
@@
As mentioned before, we have both constraints on the function $\bm{f}$ and failure regions in the input space. To this end, we define the probability of feasibility as
$g(\bm{x})=p(\mathrm{non{\text{-}}failure}|\bm{x})\ \prod_{i=1,\ldots,m}p(f_{l+i}(\bm{x})\leq 0),$
where $f_{l+1}$ to $f_{l+m}$ are the GP models of the $m$ constraint functions and $p(\mathrm{non{\text{-}}failure}|\bm{x})$ is the classification model (6). This measures the joint probability that a given point in the input space is feasible (respects all of the constraints) and is not a failure.
Using this probability, we can then define a constrained acquisition function (ECHVI):
@@
$\alpha_{\text{ECHVI}}(\bm{x})=\begin{cases}\alpha_{\text{EHVI}}(\bm{x})\cdot g(\bm{x})&\text{if }g(\bm{x})\geq\lambda,\\ 0&\text{otherwise.}\end{cases}$ (7)
which weights $\alpha_{\text{EHVI}}$ by $g$ and enforces a cut-off threshold(Gardner et al., 2014; Abdolshah et al., 2018) (here we use $\lambda=0.5$). This ensures that samples with a probability of feasibility less than $\lambda$ are excluded from consideration, while those with probabilities higher are more likely to be selected during the optimisation than those with lower probability (but still above $\lambda$). We also note that $\alpha_{\text{EHVI}}$ must be calculated with respect to the hypervolume of the feasible Pareto frontier by excluding infeasible points from the dataset.
To find the next most informative sample, we use single-objective (non-Bayesian) optimisation to find the point that maximises $\alpha_{\text{ECHVI}}$:
$\bm{x}^{*}=\underset{\bm{x}\in\mathscr{U}}{\arg\max}\ \alpha_{\text{ECHVI}}(\bm{x}).$
This is done using the L-BFGS-B(Byrd et al., 1995) algorithm which makes use of multiple restarts to avoid local maxima and avoid the discontinuity in $\alpha_{\text{ECHVI}}$. Once found, $\bm{x}^{*}$ is evaluated using $\bm{f}$ and added to the dataset $\mathscr{D}$.
III The poloidal field coil design problem
@@
III.1 The STEP baseline design
We will be working with the initial PF coil setup and limiter geometry from the baseline STEP design presented by Hudoba et al. (2023b). The design is shown in Figure 2 and the information available to us from the baseline dataset are the:
@@ III.1 The STEP baseline design
    Permissible zones for each PF coil, i.e. the region of the $RZ$-plane in which each coil can be placed without intersecting the TF coils, diagnostic systems, or other parts of the tokamak.
@@
    Limiter contour that confines the plasma equilibrium. This was constructed using the strike plate locations and made to match the geometry illustrated in Tholerus et al. (2024).
@@
    Strike plate locations, i.e. segments of the limiter in the inner and outer divertor where the legs of the plasma separatrix will strike.
@@
Figure 2: The baseline setup from Hudoba et al. (2023b) in the $RZ$-plane (due to vertical symmetry about $Z=0$ only the top half of the tokamak is shown). The separatrix (red) outlines the shape of the plasma core and the divertor legs which hit the inner (green) and outer (blue) strike plates. The initial location of the seven PF coils (grey rectangles) are displayed within their respective permissible zones (red rectangles). Note the absence of the central solenoid, which is not used in the flat-top phase of operation shown here. The isoflux constraints (black crosses) define locations at which the separatrix should pass through.
III.2 Optimisation problem
@@
III.2.1 Input space
As can be seen in Figure 2, there are seven up-down symmetric (around $Z=0$) PF coil circuits each with their own $(R^{c},Z^{c})$ centroid coordinate that is allowed to move freely such that no part of the coil leaves the permissible zone. The exceptions are the two quasi-solenoid (QS) coils, positioned above (and below) the central solenoid for magnetic shaping in the inner divertor, which are only able to move vertically. This results in a twelve dimensional input space for the optimisation problem: five pairs of $(R^{c},Z^{c})$ coordinates for the PF coils and one $Z^{c}$ coordinate for each of the two QS coils. We normalise each of the coordinates with respect to their own permissible zones so that we can work with the unit hypercube $[0,1]^{12}$ as our input space. A more detailed explanation of the normalisation process can be found in Appendix A.
III.2.2 Objectives
@@
Figure 3: The baseline setup from Figure 2, this time marked with objective and constraint quantities. Shown are the separatrix (light pink), the upper X-point (red cross), the O-point/magnetic axis (green circle), and the inner (green star) and outer (red star) strikepoints. Also shown are the flux surfaces traced out to calculate the ICL (blue plus red lines) and OCL (yellow plus green lines).
@@
The second objective is to maximise the average of the inner connection length (ICL) and the outer connection length (OCL). The connection length is the distance traced out by a helical (i.e. moving both poloidally and toroidally) magnetic field line that starts at the inner (outer) edge of the last closed flux surface (LCFS) at the midplane and ends at the inner (outer) strikepoint (Doyle et al., 2021; Albanese et al., 2023). We refer to these inner and outer midplane points as the IMP and OMP, respectively. Larger connection lengths ensure that hot plasma leaving the core edge region will travel a farther distance and therefore cool to more acceptable temperatures before hitting the strike plate. This is an important aspect of managing heat loads in the divertor region.
@@
The ICL and OCL are calculated by carrying out an integral over the length of the magnetic field line of interest (Kos et al., 2019). Tracing the magnetic field line requires the solution of a system of ordinary differential equations (ODEs) for the position vector of a point along the field line trajectory $\bm{r}(\ell)$:
$\frac{\mathrm{d}\bm{r}(\ell)}{\mathrm{d}\ell}=\bm{b}\left(\bm{r}(\ell)\right),\quad\ell\in[0,L],$ (9)
where $\bm{b}$ is a unit vector along the direction of the magnetic field $\bm{B}$ in cylindrical coordinates:
$\bm{B}\left(\bm{r}(\ell)\right)=\begin{cases}-\frac{1}{R}\frac{\partial\psi(R,Z)}{\partial Z}\\[6.0pt] \frac{F(\psi(R,Z))}{R}.\\[6.0pt] \frac{1}{R}\frac{\partial\psi(R,Z)}{\partial R}\end{cases}$
@@
Here $\psi(R,Z)$ denotes the scalar poloidal magnetic flux and $F$ the toroidal magnetic field profile. To solve this problem, we use a fourth-order Runge-Kutta method (with appropriately chosen step size $\Delta\ell$) and integrate until a terminating condition is met (see next paragraph), recording the value of $L$ obtained (i.e. the total number of steps $\Delta\ell$).
@@
In practice, however, the calculation is split into two stages. The first stage involves integrating from an initial position $\bm{r}(0)$ which is selected to be $3$mm radially outside the IMP/OMP and ending at some very small distance away from the upper X-point. The second stage integrates between the inner/outer strikepoint and the point close to the upper X-point. The length of these individual sections is then combined to return the final connection length. The two flux surfaces traced out when calculating the ICL and OCL are visualised in Figure 3. These techniques reduce the likelihood of the integrator getting stuck at the exact X-point, travelling around the LCFS (instead of going up into a divertor), and from travelling into the wrong divertor.
III.2.3 Constraints